OpenStudy (anonymous):

The number \frac{-15}{3} does NOT belong to which of the following number systems? Select one: A. Integers B. Rational Numbers C. Natural Numbers D. Complex Numbers

OpenStudy (anonymous):

@aum

OpenStudy (skullpatrol):

any ideas?

OpenStudy (anonymous):

complex numbers?

OpenStudy (skullpatrol):

nope, complex numbers have the form a +bi where a and b are real numbers, the questions asks for NOT a member of

OpenStudy (skullpatrol):

$$\Huge -\frac{15}{3}$$

OpenStudy (anonymous):

I have no idea that's why i'm asking its not integers either so

OpenStudy (skullpatrol):

that is what the question is asking for

geerky42 (geerky42):

Answer is not integer because \(-\dfrac{15}{3} = -5\), which is member of integer.

OpenStudy (anonymous):

ok then rational lol?

OpenStudy (skullpatrol):

the natural numbers are 1, 2, 3, 4, ...

geerky42 (geerky42):

Basically Rational means that number can be expressed as fraction. So that no too

OpenStudy (anonymous):

omg thanks guys

geerky42 (geerky42):

\(-5 = \dfrac{-5}{1}\) so that belong to rational too. Left us natural number. no negative number in set of natural number.

OpenStudy (anonymous):

can you help me with like 3 more?

OpenStudy (skullpatrol):

post another question please

OpenStudy (anonymous):

Which of the following demonstrates the associative property of addition? Select one: A. 2 × (5 + 3) = (2 × 5) + (2 × 3) B. 7 + (4 + 9) = (7 + 4) + 9 C. 0 × 14 = 0 D. (−8) + 8 = 0

OpenStudy (anonymous):

@geerky42 @skullpatrol

OpenStudy (anonymous):

@armine @bibby

OpenStudy (aum):

A + (B + C) = (A + B) + C is associative property of addition.